two charges one of magnitude 3 micro coulomb and the other of magnitude 5 micro coulomb if the force between them is 20 n ?


negative charge of -200 micro coulomb and a positive charge of 800 micro coulomb are separated by a distance of 30 centimeters. What is the force between the two charges?



1
Expert's answer
2022-02-20T15:58:32-0500

a)

"F=\frac{Kq_1q_2}{r^2}"


"20=\frac{9×10^9×5×3×10^{-12}}{r^2}"


"r=8.216cm"


b)


"\begin{vmatrix}
 F 
 
\end{vmatrix}=\frac{9×10^9×200×800×10^{-12}}{(30×10^{-2})^2}"



"=1.6\u00d710^4N"